If h(x)= x^2-4xy+y^2/(x+2)(y-3), find f(-1,2)
Answer by oscargut(2103)

If f(x,y)= x^2-4xy+y^2/(x+2)(y-3), find f(-1,2)
f(-1,2)= [(-1)^2-4(-1)(2)+2^2]/(-1+2)(2-3)
= (1+8+4)/(-1) = -13
Answer: -13
If you have any doubt or if you have more problems my e-mail is:
mthman@gmail.com